### v2

# Eagle: Exploring The Design Space for Multi- modal LLMs with Mixture of Encoders

Abstract
The ability to accurately interpret complex visual information is a crucial topic of multimodal large language models (MLLMs). Recent work indicates that enhanced visual perception significantly reduces hallucinations and improves performance on resolution-sensitive tasks, such as optical character recognition and document analysis. A number of recent MLLMs achieve this goal using a mixture of vision encoders. Despite their success, there is a lack of systematic comparisons and detailed ablation studies addressing critical aspects, such as expert selection and the integration of multiple vision experts. This study provides an extensive exploration of the design space for MLLMs using a mixture of vision encoders and resolutions. Our findings reveal several underlying principles common to various existing strategies, leading to a streamlined yet effective design approach. We discover that simply concatenating visual tokens from a set of complementary vision encoders is as effective as more complex mixing architectures or strategies. We additionally introduce Pre-Alignment to bridge the gap between vision-focused encoders and language tokens, enhancing model coherence. The resulting family of MLLMs, Eagle, surpasses other leading open-source models on major MLLM benchmarks.
Figure 1: Overview of Eagle. Eagle is a family of multimodal large language models (MLLMs) with a mixture of vision encoders. Left: comparisons between Eagle and existing competitive MLLMs with Vicuna-13B (Chiang et al., 2023), with Eagle achieving favorable results on all 13 benchmarks. Middle: an evolutionary road map of the design space and advanced training recipes leading to consistent and significant improvements. Right: highlights and core features of Eagle.
1 Introduction
The success of large language models (LLMs) has triggered significant interest in enabling their visual perception capability, such that they could see, understand, and reason in the real world. At the center of these multimodal large language models (MLLMs) (Fei et al., 2024) is a typical design where images are converted into a series of visual tokens by the vision encoders and appended with the text embeddings. CLIP (Radford et al., 2021) is often chosen as the vision encoder since its visual representation is aligned with the text space by pre-training on image-text pairs. Depending on the architectures, training recipes, and the way how vision tokens are injected into the language model, there exist various notable families of MLLMs such as Flamingo (Alayrac et al., 2022), BLIP (Li et al., 2022; 2023d; Dai et al., 2024), PaLI (Chen et al., 2023e), PaLM-E (Driess et al., 2023) and LLaVA (Liu et al., 2023d; c). Most of these works keep relatively low input resolutions due to the limits on pre-trained vision encoders and LLM sequence length.
Recent studies (Li et al., 2024c; Liu et al., 2024a) show that stronger vision encoder design is important for mitigating MLLM hallucinations (Liu et al., 2023a; Wu et al., 2024) and improving resolution-sensitive tasks like optical character recognition (OCR). A constellation of works thus focuses on enhancing the capability of the vision encoder. For example, scaling up the pre-training data and parameters of vision encoder (Chen et al., 2023f) or dividing images into low-resolution patches (Liu et al., 2024a; Shi et al., 2024). However, these approaches usually introduce large training resources. An efficient yet powerful strategy is to mix visual encoders pre-trained with different tasks and input resolutions, either fusing higher resolution encoders with the CLIP encoder (Luo et al., 2024; Li et al., 2024b), sequentially appending features from different encoders (Fan et al., 2024; Lin et al., 2023b; Karamcheti et al., 2024; Tong et al., 2024), or adopting more complex fusion and routing strategies to make the best of different encoders (Lee et al., 2024; Zong et al., 2024). In addition, Prismatic VLM (Liu et al., 2024b) incorporates multiple vision encoders with channel concatenation as part of their design space exploration. These “mixture-of-vision-experts” strategies are shown to be effective. However, a detailed study focusing on their designs is still lacking.
To address the above questions, our work systematically investigates the mixture-of-vision-encoders design space for improved MLLM perception. As shown in Fig. 1, our exploration of the design space consists of the following steps: 1) Benchmarking various vision encoders and searching recipes for higher resolution adaptation; 2) “Apples to apples” comparison between vision encoder fusion strategies; 3) Progressive identification of the optimal combination of multiple vision encoders; 4) Improved vision expert pre-alignment and data mixture. Our study covers the performance of vision encoders pre-trained on different tasks and resolutions (e.g., vision-language alignment (Ilharco et al., 2021; Cherti et al., 2023; Radford et al., 2021; Schuhmann et al., 2022), self-supervised learning (Oquab et al., 2023), detection (Fang et al., 2023b; a), segmentation (Kirillov et al., 2023), and OCR (Lee et al., 2023)). We use a round-robin approach to incorporate additional vision experts. Starting with the basic CLIP (Radford et al., 2021) encoder, we add one additional expert each time with the best improvement in each round.
Our work is not the first one to leverage multiple vision encoders in MLLM. However, the systematic study leads to several interesting new findings under this setting:
  • Unlocking the vision encoders during MLLM training matters. This is in sharp contrast to the LLaVA (Liu et al., 2023d; c) family and many works that consider multiple vision encoders or teachers (Lin et al., 2023b; Liu et al., 2024b; Fan et al., 2024; Kar et al., 2024; Ranzinger et al., 2024; Lee et al., 2024), where freezing the vision encoders has been a common choice.
  • Some recently proposed fusion strategies (Luo et al., 2024; Li et al., 2024b) do not show significant advantages despite their advanced designs. Instead, we find that straightforward channel concatenation stands out as a simple yet competitive fusion strategy, offering the best efficiency and performance.
  • Incorporating additional vision experts leads to consistent gain, making it a promising path to systematically enhance MLLM perception besides scaling up vision encoders. The improvement is particularly pronounced when vision encoders are unlocked.
  • We propose a pre-alignment stage where non-text-aligned vision experts are individually fine-tuned with a frozen LLM before being trained together. This stage is found to enhance the MLLM performance significantly under the mixture-of-vision-encoder design.
We finally conclude our findings into a family of MLLMs termed Eagle. Eagle is evaluated on a series of benchmarks, including visual question answering, OCR/document-related tasks, and benchmarks tailored for MLLMs. Our model attains state-of-the-art performance across different benchmarks and demonstrates obvious advantages on OCR and document understanding tasks. Using the same pre-train and supervised fine-tuning data from Cambrian-1 (Tong et al., 2024) - a concurrent family of vision-centric MLLMs sharing similar design spirits, Eagle models overall achieve better performance. We hope that the Eagle can provide a highly performant and easy-to-reproduce MLLM solution to the community.
2 Design space exploration
In this section, we show how to utilize the advantages of different vision encoders via step-by-step investigations, yielding the Eagle model family. Unlike previous methods focusing on new fusion strategies or architectures among vision encodes, our goal is to identify a set of minimalistic design to fuse different vision encoders supported with detailed ablations, removing any unnecessary parts. As shown in Fig. 2, we start by extending the basic CLIP encoder (Radford et al., 2021) to a set of vision experts with different architectures, pre-training tasks, and resolutions. With these experts, we then compare different fusion architectures and methods and study how to optimize the pre-training strategies given more encoders. We also give a detailed analysis of how to select the vision encoders to be integrated. Finally, we put all the findings together and further extend to multiple expert vision encoders with different resolutions and domain knowledge.
2.1 Base setup
We adopt LLaVA-1.5’s (Liu et al., 2023c) model architecture as the basis, which consists of a large language model (Vicuna-v1.5 7B (Chiang et al., 2023)), a vision encoder, and a projection layer. The projection layer projects the visual embedding from the vision encoder into the text embedding space.
Base training data. We adopt the same pre-training data (LLaVA-595k) as LLaVA-1.5 (Liu et al., 2023c) for the first pre-training stage, which consists of 595595595595k image text pairs. To fully examine the potential of different vision experts and fusion methods, instead of using the SFT data from LLaVA-1.5 (Liu et al., 2023c), we collect data from a series of tasks and convert them into multimodal conversations for the supervised fine-tuning (SFT) stage, denoted as Eagle1.8M in Table 1.
Table 1: Composition of the base supervised fine-tuning data (Eagle1.8M).
Total Data Size Data Source
1,809k LLaVA-1.5 (665k) (Liu et al., 2023c), DocVQA (39k) (Mathew et al., 2021), synDog-EN (50k) (Kim et al., 2022), ChartQA (28k) (Masry et al., 2022), DVQA (25k) (Kafle et al., 2018), AI2D (15k) (Kembhavi et al., 2016a), ShareGPT-4V (100k) (Chen et al., 2023b), laion-GPT4V (11k) (lai, 2023), LVIS-Instruct4V (220k) (Wang et al., 2023a), LRV-Instruct (150k) (Liu et al., 2023b), Geo170k (120k) (Gao et al., 2023), LLaVAR (20k) (Zhang et al., 2023), Visual7W (70k) (Zhu et al., 2016), Open-Hermes 2.5 (300k) (Teknium, 2023)
Base training recipes. We start from the LLaVA-1.5 (Liu et al., 2023c) recipe where the model is first pre-trained with image-text pairs for one epoch with a batch size of 256. The whole model is frozen and only the projector layer is updated in this pre-training stage. In the second stage, we further fine-tune the model on the multi-modal conversation data for one epoch with a batch size of 128. The learning rates are set to be 1e-3 for the first stage and 2e-5 for the second stage, respectively.
Base evaluation. To conduct a comprehensive comparison of various methods, we adopt 11 distinct benchmarks that span multiple tasks. These benchmarks include 1) General VQA tasks: GQA (Hudson & Manning, 2019), VizWiz (Gurari et al., 2018), MME (Fu et al., 2023), SEED (Li et al., 2023c); 2) OCR/document/chart understanding: OCRBench (Liu et al., 2023f), DocVQA (Mathew et al., 2021), ChartQA (Masry et al., 2022); 3) vision-centric tasks: POPE (Li et al., 2023e), RealWorldQA (xAI, 2024); 4) knowledge-based tasks: ScienceQA (Saikh et al., 2022), AI2D (Kembhavi et al., 2016b). To obtain an average score, we normalize each benchmark to a total score of 1,000 and then calculate the average score across all benchmarks.
Figure 2: Overview of the Eagle exploration. In this work, we explore the design space of Multi-Modal Large Language Models (MLLMs) with multiple vision encoders, aiming to identify optimized design choices and enhance MLLM perception. We collect a range of vision experts and adapt them for integration into MLLMs. A systematic comparison of popular fusion paradigms is then conducted under controlled settings. After identifying discrepancies between vision experts pre-trained on different tasks, we optimize the pre-training strategy through a pre-alignment stage and use round-robin searching to determine the optimal combination of vision encoders.
2.2 Stronger CLIP encoder
We start our exploration by upgrading the vanilla CLIP Radford et al. (2021) model since it has become a standard choice for most of the MLLMs Liu et al. (2023d; c). While CLIP models are known to benefit multimodal tasks via the text-image alignment, they also have inherent drawbacks. For instance, many existing MLLMs (Liu et al., 2023c) tend to use the pre-trained CLIP resolutions (such as $224\times 224224 × 224$ or $336\times 336336 × 336$) as their input resolutions. In these cases, the encoders often fail to capture fine-grained details that are important for resolution-sensitive tasks like OCR and document understanding (Li et al., 2024c).
To handle increased input resolution, a common practice is to use tiling where input images are divided into tiles and encoded separately (Liu et al., 2024a; Li et al., 2024c), or just interpolate the position embedding of the vision transformer model to fit high-resolution inputs (Chen et al., 2023c; d; Beyer et al., 2024). We compare these two approaches with frozen/unfrozen vision encoders under different resolutions, with the results shown in Table 2. Our findings can be summarized as follows:
  • Updating the CLIP encoder during SFT significantly improves performance at higher resolutions but slightly reduces it when using the pre-training resolution.
  • Interpolating CLIP encoder to fit the input size of $448\times 448448 × 448$ offers a strong balance between efficiency and performance, trailing the $672\times 672672 × 672$ version with less than half the tokens.
  • Despite its smaller size (0.3B vs. 5.9B) and less pre-training data, the CLIP encoder gets close with interpolation approaches InternVL’s (Chen et al., 2023f) performance under the same setting.
Table 2: Comparing different high-resolution adaption methods. “#Tok(V)” denotes the number of visual tokens. “#Parames”, “FLOPs” and “Img/Sec” denote the model size, complexity and throughput (bs=4) of the vision encoder.
Method Unfreeze Res. #Tok(V) #Params FLOPs Img/Sec Avg.
Original 336 576 0.3B 119G 197.2 616.5
Original 336 576 0.3B 119G 197.2 562.6
Interpolate 448 1024 0.3B 214G 119.5 589.7
Interpolate 448 1024 0.3B 214G 119.5 670.5
Interpolate 672 2304 0.3B 480G 56.3 674.2
Tiled-input 672 2304 0.3B 476G 51.6 673.9
InternVL 448 1024 5.9B 5669G 13.52 661.9
InternVL 448 1024 5.9B 5669G 13.52 671.5
Based on the results, we can see that Direct interpolation to $448\times 448448 × 448$ can achieve competitive performance while being more efficient. We thus use the CLIP encoder with a $448\times 448448 × 448$ input resolution while unlocking the encoder during the SFT stage.
2.3 Vision experts
To better establish the foundation for multi-vision expert fusion, we extend the toolbox with vision experts pre-trained on different tasks and resolutions, and verify our findings on high-resolution adaptation with these experts. This also helps us identify the distinct advantages of different experts. We collect a set of vision experts, including: (1) Vision-Language Alignment: CLIP (Radford et al., 2021) and ConvNeXt (Liu et al., 2022) from OpenCLIP (Ilharco et al., 2021; Schuhmann et al., 2022). (2) Object-Centric: EVA-02 (Fang et al., 2023b; a) pre-trained on detection datasets. (3) OCR: Pix2Struct (Lee et al., 2023).(4) Segmentation: SAM (Kirillov et al., 2023). (5) Self-supervised: DINOv2 (Oquab et al., 2023). We resize the output 2D feature maps of each vision encoder using bilinear interpolation or pixel shuffle (Shi et al., 2016) to ensure that the visual token number equals $10241024$.
Table 3: Comparison between different vision experts as the MLLM encoders.
Category Vision Encoder   Res. Post-process Unfreeze   Avg. Model Link
VL Alignment ConvNeXt 1024 None 654.6 ConvNeXt-XXL
682.1
Segmentation SAM 1024 Pixel Unshuffle 486.2 SAM-ViT-Large
510.5
Object Detection EVA-02 1024 Resize 543.7 EVA-02-L-Det
639.1
Text Recognition Pix2Struct 1024 Resize 598.6 Pix2Struct-02-Large
606.2
Self-Supervised DINOv2 448 None 520.7 ViT-L/14-Reg
537.3
Results in Table 3 show that unfreezing the vision experts again leads to consistent improvement, which is aligned with Sec. 2.2. In addition, results in Table 10 (see Appendix A.1) further demonstrate that MLLMs with these task-specific vision encoders achieve optimal performance in their pre-training domains. EVA-02 excels in the object hallucination evaluation benchmark POPE and general visual question answering benchmark GQA. CLIP and ConvNeXt perform well across all benchmarks, benefiting from their training on large-scale image-text pairs using contrastive loss. Conversely, while Pix2Struct excels in text recognition, it shows limited capability in object recognition and general VQA tasks, like POPE and GQA. DINOv2 and SAM, pre-trained with self-supervised learning and segmentation, struggle with text recognition tasks.
2.4 Fusion strategy
Existing MLLM frameworks have proposed various mixture-of-vision-encoder strategies, with the hope that their domain-specific strengths can be leveraged. In all cases, improvements in MLLM performance have been reported with the fusion of vision encoders. However, the roles of the fusion strategies as part of their MLLM architecture innovations, have not been decoupled and clearly studied under an “apples to apples” comparison. It is thus not entirely clear how much improvement is from the fusion strategies themselves versus the improved representations from various encoders.
We notice that existing popular fusion strategies, despite their variations in designs, can be broadly represented by the following several categories: (1) Sequence Append: directly appending the visual tokens from different backbones as a longer sequence (Fan et al., 2024; Kar et al., 2024); (2) Channel Concatenation: concatenating the visual tokens along the channel dimension without increasing the sequence length (Lin et al., 2023b; Karamcheti et al., 2024); (3) LLaVA-HR: injecting high-resolution features into low-resolution vision encoders using mixture-of-resolution adapter (Luo et al., 2024); (4) Mini-Gemini: using the CLIP tokens as the low-resolution queries to cross-attend another high-resolution vision encoder in the co-located local windows (Li et al., 2024b). (5) Deformable Attention: a new baseline we introduce on top of Mini-Gemini, where the vanilla window attention is replaced with deformable attention (Zhu et al., 2021). Fig. 2 gives a detailed illustration of these fusion strategies. To better study them, we choose “CLIP+ConvNeXt” and “CLIP+ConvNeXt+SAM” as the base multi-encoder combinations to perform comparisons.
Table 4: Comparison of different fusion methods for different vision experts. “#Token(V)” denotes the number of visual tokens. “#Tokens/s” denotes the inference throughtput of the whole pipeline.
Vision Encoders Fusion #Token(V) #Tokens/s #Params Avg.
CLIP + ConvNeXt Seq. Append 2048 46.1 1200M 690.5
Channel Concat. 1024 47.3 1184M 681.5
LLaVA-HR 1024 47.0 1219M 678.7
Mini-Gemini 1024 45.3 1201M 672.5
Deformable Attn. 1024 47.3 1201M 674.3
CLIP + ConvNeXt + SAM Seq. Append 3072 40.3 1529M 686.2
Channel Concat. 1024 46.3 1495M 690.4
Our study in Table 4 shows that Channel Concatenation stands out with the best performance, expandability, and efficiency. The “injection-based” methods, such as LLaVA-HR, Mini-Gemini and Deformable Attention, are in general less competitive on TextVQA (Singh et al., 2019) and OCRBench (Liu et al., 2023f), performing worse than using ConvNeXt alone as the vision encoder. Although sequence append shows comparable performance to channel concatenation, it faces the challenge of handling significantly increased sequence lengths with additional vision encoders.
2.5 Vison-language Pre-Alignment
Figure 3: The proposed training strategy of Eagle. It consists of three progressive stages, including vision-language pre-alignment training, joint-project training and supervised fine-tuning. These stages effectively leverage public data from diverse sources, ranging from noisy image-text pairs on the web to high-quality caption, VQA, and multimodal dialogue datasets.
As shown in Table 3, encoders pre-trained exclusively on vision tasks (e.g., detection, OCR, and segmentation) are less competitive compared to those pre-trained on vision language alignment. This is possibly due to representational inconsistencies when integrated with large language models. Additionally, when combining different encoders, there is a gap between these encoders, creating difficulties in the training process. To address this feature inconsistency, we propose a Pre-Alignment training stage that first aligns each individual vision encoder with the same large language model, fostering better synergy between visual and linguistic capabilities.
Fig. 3 depicts our pre-alignment strategy. Instead of training a projector to simultaneously align multiple vision experts as in LLaVA’s (Liu et al., 2023c) original pre-training strategy, we first align the representation of each individual expert with a smaller language model (Vicuna-7B in practice) using next-token-prediction supervision. As shown in Fig. 3, with pre-alignment, the whole training process consists of three steps: 1) training each pre-trained vision expert with their own projector, while keeping the language model frozen; 2) combining all vision experts from the first step and training both the projector and vision experts; 3) training the whole model on SFT data.
Table 5: The effectiveness of Pre-alignment.
CLIP Vision Expert (X) Unfreeze Pre-align Avg.
CLIP-448 SAM-1024 630.6
648.5
662.5
672.3
CLIP-448 ConvNext-1024 652.0
670.1
681.5
686.2
CLIP-448 Pix2Struct-1024 653.5
665.7
673.7
680.4
CLIP-448 EVA-02-L-1024 630.2
645.2
659.2
668.2
To verify the proposed method, we compare the pre-alignment strategy with the normal two-stage training strategy in Table 5, considering both freezing and unfreezing vision experts for comparison. As shown in Table 5, although unfreezing the vision experts during SFT helps improve performance by updating the vision experts to fit the language model, the Pre-Align strategy more effectively mitigates the inherent biases of each vision expert and stabilizes the training process, subsequently improving overall performance.
2.6 Extension to multi-experts
With the optimized strategies and training recipes of incorporating individual vision experts, we consider the incorporation of even more vision experts to push the limit. To conduct the search in a systematic manner, we adopt a step-by-step greedy strategy to incorporate additional vision experts.
We consider the vision experts discussed in Section 2.3 for experiments. We mark CLIP, ConvNeXt, SAM, DINOv2, Pix2Struct, and EVA-02-L as CL, CN, SA, DI, PS, and EV, respectively. A round-robin scheme, as shown in Fig. 4, is adopted. We first use the two top-performing vision encoders, CLIP and ConvNeXt, as the basis and gradually add one more vision encoder each time. In each round, the best-performing vision encoder combination is retained for the next round.
#Encoder Encoder Combination Config #Params (M) FLOPs (G) Img/Sec Avg.
2 CL + CN X2 1155.2 3347.2 18.1 681.5
3 CL + CN + DI 1460.6 3659.9 15.1 685.4
CL + CN + SA 1463.9 4657.8 8.8 690.4
CL + CN + PS 1669.6 4373.2 6.9 685.1
CL + CN + EV X3 1459.6 4280.9 9.1 690.7
4 CL + CN + EV + DI 1765.1 4593.6 8.3 688.0
CL + CN + EV + SA 1768.4 5591.5 5.9 689.4
CL + CN + EV + PS X4 1974.1 5306.9 5.0 694.6
5 CL + CN + EV + PS + DI 2279.5 5619.5 4.7 684.7
CL + CN + EV + PS + SA X5 2282.8 6617.4 3.8 697.1
6 CL + CN + EV + PS + SA + DI X6 2588.2 6930.1 3.6 686.8
Figure 4: Results of vision expert selection process. CL, CN, EV, PS, SA and DI denote CLIP, ConvNeXt, EVA-02, Pix2Struct, SAM and DINOv2, respectively. (Left) The performance of various vision encoder combinations is presented, highlighting how different pairings influence overall effectiveness. “#Parames”, “FLOPs” and “Img/Sec” denote the model size, complexity and throughput (bs=4) of the vision encoder. (Right) The curve illustrates the average score as the number of vision encoders increases. Each point on the curve represents the best-performing combination for the corresponding number of vision encoders. No Pre-Alignment is used in this comparison.
Fig. 4 reveals several insights. Generally, introducing additional vision encoders enhances the performance. This indicates that the distinct advantages of different encoders can be preserved and utilized; for example, integrating the EVA-02 encoder improves metrics on the POPE benchmark. Although individual metrics may vary, the aggregated performance shows an upward trend, as evidenced by normalized average metrics, suggesting that the overall efficacy of the system is enhanced with more encoders. Also, Fig. 4 shows that the best combination of vision experts are CLIP, ConvNeXt, SAM, Pix2Struct, and EVA-02. We will use this recipe in the final model.
3 Experiments
In this section, we take the findings and the best-explored design from Section 2 and compare them against the current state-of-the-art MLLMs on different tasks.
3.1 Implementation details
Language models. We use Vicuna-v1.5-7B (Chiang et al., 2023), Llama3-8B (AI@Meta, 2024) and Vicuna-v1.5-13B (Chiang et al., 2023) as the language models.
Vision encoders. We follow the best X4 and X5 configurations, where the interpolated CLIP-448 and pre-aligned vision experts are channel-concatenated, and trained following the exact best training recipes in Figure 3 and Table 5.
Training recipe. On Eagle1.8M, we follow the base recipe in Section 2.1 with encoder learning rate the same as SFT (2e-5). On Cambrian data, we follow Tong et al. (2024) with PT/SFT bs=1024.
3.2 Main results
Evaluation on visual question answering tasks. We compare Eagle model series across three Visual Question Answering (VQA) benchmarks, including GQA (Hudson & Manning, 2019), VQAv2 (Goyal et al., 2017) and VizWiz (Gurari et al., 2018). As shown in Table 6, Eagle-X5 achieves state-of-the-art performance on GQA and VQAv2, underscoring the advantages with additional vision experts.
Table 6: Main results with base training data. SQAI denotes image split of ScienceQA.
Model
MME
MMB
SEED
MathVista
MMMU
POPE
SQAI
GQA
VizWiz
VQAv2
OCR
TextVQA
ChartQA
Vicuna-7B & Qwen-7B LLaVA-1.5 (Liu et al., 2023c) 1510 64.3 58.6 - - 85.9 66.8 62.0* 50.0 78.5* 297 58.2 -
LLaVA-NeXt (Liu et al., 2024a) 1519 67.4 70.2 34.6 35.8 86.5 70.1 64.2* 57.6 80.0* 490 64.9 -
InternVL (Chen et al., 2023f) 1525 - 65.4 - - 86.4 - 62.9* 52.5 79.3* - 57.0 -
LLaVA-HR (Luo et al., 2024) 1554 - 64.2 - - 87.6 65.1 64.2* 48.7 81.9* - 67.1 -
Monkey (Li et al., 2024c) - - - - - - - 60.7* 61.2* 80.3* 514 67.6 65.1
Mini-Gemini (Li et al., 2024b) 1523 65.8 - 32.2 36.8 - 71.1 64.5* - - 477 65.2 -
Eagle-X5 1528 68.4 73.9 37.0 36.3 88.8 70.0 64.9* 54.4 83.4* 529 71.2 67.7
Eagle-X5 (+Pre-Align) 1582 69.7 73.7 38.2 38.0 88.7 71.9 64.6* 58.7 83.6* 566 71.9 69.3
Vicuna-13B LLaVA-1.5 (Liu et al., 2023c) 1531 67.7 61.6 - 36.4 85.9 71.6 63.3* 53.6 80.0* 331 61.3 -
LLaVA-NeXt (Liu et al., 2024a) 1575 70.0 71.9 35.3 36.2 86.2 73.5 65.4* 60.5 82.8* 514 67.1 62.2
InternVL (Chen et al., 2023f) 1546 - - - - 87.1 - 63.9* 54.6 80.2* 517 58.7 -
LLaVA-UHD (Xu et al., 2024) 1535 68.0 - - - 89.1 72.0 65.2* 56.1 81.7* - 67.7 -
LLaVA-HR (Luo et al., 2024) 1540 - 64.5 - - 87.8 68.1 64.8* 57.9 82.6* - 68.1 -
Mini-Gemini (Li et al., 2024b) 1565 68.6 70.6 37.0 37.3 - 71.9 65.8* - - 466 65.9 56.6
Eagle-X5 1609 69.2 74.1 38.8 36.6 87.8 72.7 66.2* 59.3 83.8* 574 74.2 69.9
Eagle-X5 (+Pre-Align) 1605 71.6 74.9 42.7 38.5 89.2 75.5 64.6* 60.9 84.5* 598 73.3 72.1
Evaluation on OCR and chart understanding tasks. To evaluate the OCR, document, and chart understanding capabilities of Eagle, we benchmark our model on OCRBench (Liu et al., 2023f), TextVQA (Singh et al., 2019), and ChartQA (Masry et al., 2022). As illustrated in Table 6, our model significantly surpasses competitors on TextVQA, benefiting from its high-resolution architecture and integration of different vision encoders. Notably, Eagle maintains a straightforward design, supporting up to 1024x1024 resolution without requiring complex tile decomposition of images.
Fig. 5 shows some examples of OCR and document understanding cases. With high-resolution adaptation and more vision experts, our model can identify small text within images and accurately extract information according to the users’ instructions. To better understand the benefits of introducing experts pre-trained on other vision tasks, we visualize the results of a model with only the ConvNeXt and CLIP vision encoders, compared to the results of Eagle-X5 in Fig. 5. With the full set of vision encoders, the model can successfully correct mistakes, showing that even when equipped with high-resolution vision encoders pre-trained on vision-language alignment, the model’s abilities can still be enhanced by integrating additional vision experts pre-trained on diverse vision tasks.
Evaluation on multimodal benchmarks. We evaluate Eagle on seven benchmarks for MLLMs to demonstrate its capabilities from different perspectives, including MME (Fu et al., 2023), MMBench (Liu et al., 2023e), SEED (Li et al., 2023c), MathVista (Lu et al., 2024), MMMU (Yue et al., 2024), ScienceQA (Saikh et al., 2022), and POPE (Li et al., 2023e). Specifically, MME, MMBench, and SEED assess the overall performance on various real-world tasks based on reasoning, recognition, knowledge, and OCR. MMMU focuses on challenging problems from diverse domains that require college-level knowledge. POPE evaluates the visual hallucinations of MLLMs. The metrics used in our paper adhere to the default settings of these benchmarks. We report the perception score for MME, the en_dev split for MMBench, the image split of SEED, the test-mini split of MathVista, the val split of MMMU, the F1-score of POPE, and the image split of SQA to align with the reported scores from other models.
Table 7: Comparison between different training strategies. “1 epoch” means we train Eagle for 1 epoch in the supervised fine-tuning stage. ‘unlock*” means we unlock vision encoders in the pre-training stage. More details in Table. 13.
Config Summary Pre-align Pre-train Fine-tune Avg.
1 epoch LLaVA-595K Eagle1.8M 697.1
2 epoch LLaVA-595K Eagle1.8M 698.3
1 epoch, unlock* LLaVA-595K Eagle1.8M 698.0
1 epoch, unlock* LLaVA-595K+Eagle1.8M Eagle1.8M 699.5
1 epoch Eagle1.8M LLaVA-595K Eagle1.8M 706.6
1 epoch, unlock* Eagle1.8M LLaVA-595K Eagle1.8M 707.1
1 epoch, unlock* LLaVA-595K+Eagle1.8M LLaVA-595K Eagle1.8M 707.8
1 epoch, unlock* LLaVA-595K+Eagle1.8M LLaVA-595K+Eagle1.8M Eagle1.8M 708.9
From the data presented in Table 6, Eagle consistently surpasses existing models across various MLLMs on SEED and MME, demonstrating the comprehensive knowledge and reasoning abilities of Eagle. With the help of vision encoders on object-centric tasks, Eagle also achieves the best performance on the POPE benchmark. Additionally, the Pre-Alignment strategy discussed in Sec. 2.5 has been found to further enhance performance when integrating multiple task-specific vision backbones. This approach not only mitigates the inherent biases of each vision expert and the synergy between different modalities but also establishes a robust framework for multiple-expert fusion.
Study on more advanced training recipes. Table. 7 presents our step-by-step experiments to study the training recipes. We found that the best recipe is to first pre-align each vision expert on LLaVA-595K + Eagle1.8M. In the pretraining stage, we combine all vision experts from the first step and training both the projector and vision experts on LLaVA-595K + Eagle1.8M. Finally, we train the whole model on the Eagle1.8M.
Comparison with Cambrian-1. Using the same pre-training and supervised fine-tuning datasets from Cambrian-1 (Tong et al., 2024), Eagle demonstrates superior performance across all the evaluated benchmarks without bells and whistles. As shown in Table 8, Eagle outperforms the Cambrian-1 counterparts considerably for the OCR and Chart category. Consistent improvements are also observed for the General, Knowledge, and Vision-Centric categories, showing the robustness and generalization ability of the improved perception design in Eagle.
Table 8: Results using the same training data as Cambrian-1 (Tong et al., 2024). SQAI denotes ScienceQA-IMG (Saikh et al., 2022). RWQA denotes the RealworldQA (xAI, 2024).
Model Knowledge General OCR and Chart Vision-Centric
Avg
SQAI
MMMU
MathVista
AI2D
Avg
MME
MMB
SEED
GQA
Avg
ChartQA
OCR
TextVQA
DocVQA
Avg
MMVP
RWQA
Llama3-8B
MGM-HD 55.7 75.1 37.3 37.0 73.5 72.7 1606 72.7 73.2 64.5 62.9 59.1 47.7 70.2 74.6 40.4 18.7 62.1
Cambrian-1 61.3 80.4 42.7 49.0 73.0 73.1 1547 75.9 74.7 64.6 71.3 73.3 62.4 71.7 77.8 57.6 51.3 64.2
Eagle-X5 65.2 84.1 43.5 56.9 76.2 74.0 1587 75.5 76.5 64.9 77.0 80.7 62.6 76.7 87.1 59.6 52.0 67.2
Vicuna-13B
MGM-HD 54.1 71.9 37.3 37.0 70.1 70.7 1597 68.6 70.6 63.7 60.8 56.6 46.6 70.2 69.8 38.4 19.3 57.5
Cambrian-1 60.2 79.3 40.0 48.0 73.6 73.7 1610 75.7 74.4 64.3 71.3 73.8 61.9 72.8 76.8 52.2 41.3 63.0
Eagle-X5 63.8 82.6 42.2 54.6 73.8 74.6 1651 75.7 75.0 65.0 75.7 78.6 62.4 74.9 86.7 54.8 44.6 65.0
Yi-34B
MGM-HD 62.4 77.7 48.0 43.4 80.5 76.2 1659 80.6 75.3 65.8 68.1 67.6 51.8 74.1 78.9 52.3 37.3 67.2
Cambrian-1 67.0 85.6 49.7 53.2 79.7 76.8 1689 81.4 75.3 65.8 71.9 75.6 60.0 76.7 75.5 60.3 52.7 67.8
Eagle-X5 68.6 85.5 53.2 57.9 79.1 76.3 1677 81.0 75.6 64.9 75.4 77.2 62.4 78.8 83.0 59.8 50.0 69.5
4 Related work
Multimodal large language models. Our work is related to the general architecture design of multimodal large language models. Besides the line of representative open-source research mentioned in the introduction section, other notable families of MLLMs include, but are not limited to MiniGPT-4 (Zhu et al., 2023; Chen et al., 2023a), Lynx (Zeng et al., 2023), Otter (Li et al., 2023b; a), Qwen-VL (Bai et al., 2023), CogVLM (Wang et al., 2023b; Hong et al., 2024), VILA (Lin et al., 2023a), GPT-4V (Achiam et al., 2023), Gemini (Team et al., 2023), and Llama 3.1 (Dubey et al., 2024). Depending on how vision signals are integrated into the language model, MLLMs can be broadly categorized into “cross-modal attention” ones and “prefix-tuning” ones (Yin et al., 2024). The former injects visual information into different layers of LLMs using cross-modal attention (Alayrac et al., 2022; Li et al., 2023b), whereas the latter views the visual tokens as part of the language token sequence and directly append them with text embeddings (Liu et al., 2023d; Li et al., 2022; 2023d). Our model belongs to the prefix-tuning family by following a LLaVA-styled multimodal architecture. Considering that MLLM is a fast-growing field, readers are recommended to refer to more detailed studies and surveys such as Yin et al. (2024); Li et al. (2024a); Fei et al. (2024).
Vision encoder design for MLLMs. Our work is closely related to research focusing on improved vision encoder designs for MLLM. Early works (Liu et al., 2023d; Li et al., 2022) usually adopted vision encoder pre-trained on vision-language alignment tasks such as CLIP (Radford et al., 2021) and EVA-CLIP (Sun et al., 2023). Stronger vision encoders such as SigLIP (Zhai et al., 2023) and InternVL (Chen et al., 2023f) have been proposed to improve vision-language tasks with improved designs, larger model sizes, and better training recipes. Considering that the models are often pre-trained on low-resolution images and lack the ability to encode fine-grained details, higher resolution adaptation is often performed to increase the MLLM input resolution (Chen et al., 2023c; d; Beyer et al., 2024; Chen et al., 2024).
Besides higher resolution adaptation, models such as LLaVA-NeXT (Liu et al., 2024a), LLaVA-UHD (Xu et al., 2024), Monkey (Li et al., 2024c), InternLM-XComposer (Dong et al., 2024), and InternVL (Chen et al., 2024) use adaptive tiling to divide high-resolution input into lower resolution patches that are processed separately. Although the ability to handle higher resolution is similarly made possible with the introduction of additional vision experts, its spirit is slightly orthogonal to the tiling techniques in terms of diversifying the visual representations. Both techniques are compatible and can be combined together.
Our work is most related to existing models using multiple vision encoders for improved perception. Mini-Gemini (Li et al., 2024b) and LLaVA-HR (Luo et al., 2024) propose to fuse high-resolution visual features into the low-resolution visual tokens. Apart from the resolution issue, these pre-trained vision encoders may lack specific abilities such as reading text and localizing objects. Hence, a series of works have integrated vision models pre-trained on different tasks for more comprehensive capabilities. For example, Mousi (Fan et al., 2024), Prismatic VLM (Karamcheti et al., 2024) and Brave (Kar et al., 2024) fuse visual tokens from different vision encoders by concatenating along the channel or token direction. There are also more complex approaches, including knowledge distillation Ranzinger et al. (2024), augmenting the input prompt with the information output by the vision experts (Lee et al., 2024; He et al., 2024; Liu et al., 2024b), or using a routing network to assign input to proper vision experts (Zong et al., 2024). In particular, Prismatic VLM (Karamcheti et al., 2024) systematically explores the design space of MLLMs across various dimensions, including data, training recipes, and notably. It also includes the vision encoder ensemble as part of its enhancement strategies. However, it lacks a comprehensive ablation study and discussion addressing the challenges associated with combining multiple vision encoders.
Figure 5: Qualitative comparison of different numbers of vision experts. Baseline means Eagle with only CLIP+ConvNext. More Vision Experts denotes the Eagle-X5 model. We highlight a failure case in RED. BLUE indicates the correct answers. With more vision experts, Eagle can more precisely capture the information in the image and provide the correct answer.
5 Conclusion
We conduct an in-depth analysis study on the design space for integrating vision encoders for multimodal large language models. Unlike previous works that focus on designing novel fusion paradigms, we find systematic design choice matters and discover a series of useful techniques. Step by step, we optimize the training recipes of individual vision encoders, identify an extendable and efficient fusion method, and gradually combine vision encoders with different domain knowledge. The results show the importance of basic design space. We hope our work can serve as a new basis and bring new inspiration for the vision encoder design for MLLM.
6 Acknowledgments
The team would like to thank Yunhao Fang and Jason Lu for sharing the data and training recipes, and Yawen Luo for the assistance on figure editing. We thank Wei Ping, Zhuolin Yang, Wenliang Dai, Nayeon Lee, Boxin Wang, Ilia Karmanov, Lukas Voegtle, Philipp Fischer, Matthieu Le and Tuomas Rintamaki for their assistance on the internal codebases. We also thank the valuable discussions and input from Zhiqi Li, Guo Chen, Shilong Liu, Jihao Liu, Ming-Chang Chiu, Yunze Man, Shiyi Lan, Nadine Chang, Maying Shen, Vibashan VS, Jenny Schmalfuss, Jose Alvarez, Amala Sanjay Deshmukh, Mike Ranzinger, Greg Heinrich, Pavlo Molchanov, Vidya Murali, Parthasarathy Sriram, Mohammad Shoeybi, Song Han, Ofri Masad, Osvald Nitski, Qing Miao, Yao Xu, Jane Scowcroft, Dmitry Chichkov and Padmavathy Subramanian. Finally, the team would like to thank the Hugging Face Team that for the support of ZERO GPU demo, and the NVIDIA infrastructure team for their prompt and helpful assistance. Min Shi is partly supported by NSF Award #2427478 and #2229873.
Appendix A Appendix
A.1 Benchmark details
In this section, we provide the additional benchmark details for the tables in Section 2. The detailed comparison of different adaptation methods for the CLIP encoder are shown in Table 9. Table 10 shows the comparison between different vision encoders on all the adopted benchmarks. Table 11 list the detailed results on the vision encoder fusion methods. Table 12 shows the comparison between different vision encoder combinations on each benchmark.
Table 9: Comparison of different high-resolution adaption methods to strengthen CLIP model (336x336). RWQA denotes the RealworldQA (xAI, 2024).
Vision Encoder Unfreeze Res GQA VizWiz MME SEED OCR DocVQA ChartQA AI2D POPE RWQA SQA Avg.
Original 336 63.2 55.1 1574 70.2 354 59.6 42.1 71.1 86.7 58.0 72.5 616.5
Original 336 60.9 54.0 1501 61.8 305 56.4 24.5 69.3 80.9 53.6 72.0 562.6
Interpolate 448 62.7 53.8 1470 69.1 299 58.4 37.7 71.2 85.7 55.3 69.6 589.7
Interpolate 448 65.6 57.8 1534 73.7 526 65.0 61.1 73.7 87.3 57.7 71.5 670.5
Interpolate 672 64.9 55.7 1503 72.4 509 64.6 62.0 72.2 87.1 57.4 71.2 674.2
Tiled-input 672 63.0 54.9 1529 72.5 435 64.9 65.7 71.5 87.6 57.0 71.4 673.9
InternVL 448 63.6 56.9 1537 71.7 529 65.0 58.6 72.9 87.4 59.2 70.2 661.9
InternVL 448 65.6 57.8 1534 73.7 526 65.0 61.1 73.7 87.3 58.8 71.5 671.5
Table 10: Comparison between different vision experts as the MLLM encoders.
Category Vision Encoder Unfreeze Res GQA VizWiz MME SEED OCR DocVQA ChartQA AI2D POPE RWQA SQA Avg.
VL Alignment ConvNeXt 1024 63.3 53.5 1526 70.6 404 70.4 60.8 71.6 87.5 57.1 68.6 635.0
1024 63.3 54.4 1510 72.9 518 77.9 67.0 72.1 88.1 58.8 68.6 659.7
Segmentation SAM 1024 57.3 49.0 1216 56.9 38 20.1 17.4 69.2 84.3 49.2 66.8 471.1
1024 60.2 51.5 1291 65.9 35 21.2 17.8 70.7 86.4 54.1 65.7 494.7
Object Detection EVA-02 1024 63.1 51.1 1359 69.2 123 25.6 25.2 71.2 88.5 57.9 66.1 523.5
1024 64.3 55.5 1449 72.7 358 57.1 57.5 72.2 88.3 59.6 67.7 614.4
Text Recognition Pix2Struct 1024 53.1 48.1 1296 53.4 460 71.0 61.0 69.6 79.2 46.7 65.5 578.7
1024 54.9 47.3 1262 55.1 472 72.5 62.0 68.7 80.0 49.3 66.6 584.8
Self-Supervised DINOv2 448 62.4 53.1 1438 67.4 41 20.2 17.3 70.7 85.3 53.3 67.1 503.1
448 64.2 55.5 1466 71.8 45 20.4 17.5 71.4 87.4 57.3 67.8 518.0
Table 11: Comparison of different fusion methods for different vision experts. “#Tokens(V)” denotes the number of visual tokens.
Vision Encoders Fusion #Tokens(V) GQA VizWiz MME SEED OCR DocVQA ChartQA AI2D POPE RWQA SQA Avg.
CLIP + ConvNeXt Seq. Append 2048 64.8 54.5 1563 73.4 532 77.7 67.6 72.4 87.9 61.2 68.8 690.5
Channel Concat. 1024 63.2 48.0 1497 73.5 551 77.7 67.0 72.4 88.3 59.1 70.7 681.5
LLaVA-HR 1024 64.5 57.2 1538 72.0 498 74.5 63.8 72.3 87.7 59.2 68.7 678.7
Mini-Gemini 1024 65.3 56.9 1548 72.9 478 68.3 63.2 71.5 87.3 59.7 69.4 672.5
Deformable Attn. 1024 64.0 57.3 1504 72.7 463 69.5 64.4 73.3 87.4 62.8 68.9 674.3
CLIP + ConvNeXt + SAM Seq. Append 3072 64.3 53.6 1539 73.2 525 77.9 67.0 72.3 87.4 60.1 69.5 686.2
Channel Concat. 1024 63.3 55.9 1528 73.3 545 78.9 67.2 72.3 88.4 59.2 70.0 690.4
Table 12: Detailed comparison on vision encoder combinations.
#Encoder Encoder Combination GQA VizWiz MME SEED OCR DocVQA ChartQA AI2D POPE RWQA SQA Avg.
2 CL + CN 63.2 48.0 1497.0 73.5 551.0 77.7 67.0 72.4 88.3 59.1 70.7 681.5
3 CL + CN + DI 63.3 55.9 1528.0 73.3 545.0 78.9 67.2 72.3 88.4 59.2 70.0 690.4
CL + CN + SA 64.6 55.3 1504.0 73.3 526.0 75.7 64.9 72.1 88.3 61.1 70.9 685.4
CL + CN + PS 63.2 51.4 1497.0 73.3 550.0 78.5 65.9 73.1 87.7 60.3 70.5 685.1
CL + CN + EV 63.2 51.7 1565.0 73.9 538.0 77.7 67.8 73.6 89.0 61.4 69.4 690.7
CL + CN + EV + DI 63.6 54.9 1512.0 73.8 547.0 77.0 66.7 73.1 88.9 60.4 69.7 689.4
CL + CN + EV + SA 64.3 57.7 1533.0 73.7 521.0 75.2 65.3 72.2 88.5 61.1 70.2 688.0
CL + CN + EV + PS 64.8 56.5 1561.0 73.4 540.0 78.8 67.5 72.2 88.4 59.9 70.5 694.6
5 CL + CN + EV + PS + SA 64.7 59.1 1528.0 73.9 529.0 78.6 67.8 72.9 88.8 62.2 69.7 697.1
CL + CN + EV + PS + DI 64.7 54.1 1506.0 73.7 541.0 75.1 64.9 72.7 88.3 60.0 70.3 684.7
6 CL + CN + EV + PS + SA + DI 63.8 57.8 1512.0 73.5 525.0 75.1 65.8 71.8 88.4 61.4 69.9 686.8
Table 13: Comparison between different training strategies. ‘1 epoch” means we train Eagle for 1 epoch in the supervised fine-tuning stage. ‘unlock*” means we unlock vision encoders in the pre-training stage.
Config Prealign Pretrain Finetune
GQA
MME
OCR
SciQA
POPE
DocVQA
ChartQA
SEED
Vizwiz
AI2D
RWQA
Avg.
1 epoch llava595k Eagle1.8M 64.7 1528 52.9 69.7 88.8 78.6 67.7 73.9 59.1 72.8 62.2 697.1
2 epoch llava595k Eagle1.8M 65.4 1539 51.4 70.3 87.9 79.8 67.9 73.8 58.5 73.5 62.7 698.3
1 epoch, unlock* llava595k Eagle1.8M 64.1 1541 54.4 71.5 88.5 79.1 68.5 74.0 56.6 72.2 61.9 698.0
1 epoch, unlock* llava595k+Eagle1.8M Eagle1.8M 65.3 1545 54.8 70.5 88.5 78.8 68.4 73.5 57.5 72.0 62.9 699.5
1 epoch Eagle1.8M llava595k Eagle1.8M 64.5 1579 55.4 71.6 88.7 79.3 68.9 73.9 58.9 73.9 63.4 706.6
1 epoch, unlock* Eagle1.8M llava595k Eagle1.8M 65.1 1561 56.0 71.3 88.5 79.3 69.5 74.2 59 73.5 63.4 707.1
1 epoch, unlock* llava595k+Eagle1.8M llava595k Eagle1.8M 64.5 1570 56.0 71.7 89 79.7 68.9 74.0 58.5 74.0 63.7 707.8
1 epoch, unlock* llava595k+Eagle1.8M llava595k+Eagle1.8M Eagle1.8M 64.6 1582 56.6 71.9 88.7 79.8 69.3 73.7 58.7 73.5 63.9 708.9
A.2 Additional qualitative results
Fig. 6 demonstrates that our model can perform reasoning based on the knowledge to answer some difficult questions. Fig. 7 shows some examples of OCR and document understanding cases.
Figure 6: Qualitative results of Eagle on VQA tasks requiring college-level knowledge or deliberate reasoning.
Figure 7: Qualitative samples on OCR and document understanding tasks. Eagle is able to extract useful information from small text.